A test is worth 200 points. The test consists of true and false questions worth 6 points each and multiple choice questions are worth 13 points each. How many multiple choice questions are on the test and true and false question?
Answer by stanbon(75887)

A test is worth 200 points. The test consists of true and false questions worth 6 points each and multiple choice questions are worth 13 points each. How many multiple choice questions are on the test and true and false question?
----
6t + 13m = 200
t = (-13/6)m + 200/6
Graph with your calculator and look for
points with 2 natural numbers using the
"table" button on a TI calculator.
-------------------------
Multiple = 2 ; t/f = 29
OR
Multiple = 8 ; t/f = 16
Or
Multiple = 15 ; t/f = 3
----
Cheers,
Stan H.